OpenStudy (anonymous):

I will medal. (: What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 A. {(31/8, 25/8)} B. {(25/8, 31/8)} C. {(45/2, 31/2)} D. {(31/2, 45/2)}

OpenStudy (anonymous):

Hello @brooke_rogers1 my name is Cylah and I can help you with this equation. Could you please tell me what grade you are in

OpenStudy (anonymous):

I am a junior but I'm studying for ACT so i have to go back over all my math that i've learned.

OpenStudy (anonymous):

Ok, well what would be your best guess. I will explain the answer afterwards.

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

A would be correct, good try but the reason A is correct is because 5x-3y-10=0 => y=(5/3)x-10/3 x+y-7=0 => y=-x+7 {y=(5/3)x-10/3 {y=-x+7 multiply second equation in (-1) => (-1)(y=-x+7) so y=(5/3)x-10/3 -y=x-7 add two equation (add left side with left side & rught side with right side) then: 0=(8/3)x-31/3 (8/3)x=31/3 x=31/8 y=-x+7 => (x=31/8)=>y=31/8 =>y=25/8 so : x=31/8 & y=25/8

OpenStudy (anonymous):

Is that easy to understand?

OpenStudy (anonymous):

Yeah..i think i actually got it, thank you so much, i really appreicate your help!

OpenStudy (anonymous):

Thanks for asking. Here anytime for help